• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 15:23:04 on March 8,2002

    Entry number 1201

    keyword=status for E98-108 running

    Detectors (aside from RICH) are installed and wired up. Right arm HV controls complains about a missing card; Javier Gomez is helping with it. The information on the PbG HV values now must be re-entered by hand. At the moment we only have the values from last April 2001 to enter.

    RICH is installed and cabling is being completed. Low voltage is being applied.

    Target is cooled, both 4 cm and 15 cm cells are available. No cryomotion survey was done -- we NEED to get the survey done at the end of the experiment.

    John LeRose is ramping the right arm magnets (and probably the left arm as well). We should be able to start cosmic runs in the next few hours.